Gucchi:

The function f(x) = 3,267(1 + 0.02)x represents the amount of money in a savings account where x represents time in years. What does the 3,267 represent? The amount of money in the savings account after one year The amount of money added to the savings account each year The initial amount of money placed in the savings account The growth rate

Gucchi:

Am i right with C?...

Gucchi:

@axie

axie:

And why would that be? (prove answer)

Gucchi:

wait actually, c?

Gucchi:

3,267 represents The initial amount of money placed in the savings account

axie:

f(x) = 3,267(1 + 0.02)^x This is in the form y = a b^x where a is the initial amount b is the growth rate x is the time 3267 is the initial amount 1.02 is the growth rate, so it grows by .02 or 2 percent f(x) = 3,267(1 + 0.02)^x This is in the form y = a b^x where a is the initial amount b is the growth rate x is the time 3267 is the initial amount 1.02 is the growth rate, so it grows by .02 or 2 percent and x is the time So what does a=?

Gucchi:

initial amount?

axie:

Yuh!

Gucchi:

so

Gucchi:

it is c

axie:

Yup!

Gucchi:

ty